Find the probability of getting a three first and then a four when the die is rolled twice.: Find the probability of getting a three first and then a four when the die is rolled twice.
Answer by Fombitz(1755)

P(3)=1/6
p(4)=1/6
Since they're independent events, the probabilities multiply.
P(3 then 4)=1/6*1/6=1/36